Actions
RunPlan20150719¶
July | 19 | ||
estimated time | 9:30-15:00 | 15:00-20:00 | 20:00-00:00 |
MOM | R. Bayes | R. Bayes | R. Bayes |
Lead Shifter | Jan | Celeste | Paolo |
Shifter 1 | Tanaz | Pavel | Dejan |
BLOC | Celeste | Celeste | Joe |
SOC | |||
Experimenter | |||
Aim | Magnet Commissioning | Magnet Commissioning | Tracker Timing Studies |
Beam? | No | No | Yes |
Additional Personnel | Ed Overton | ||
Paul Rubinov |
Shift plans¶
Magnet Commissioning¶
- To be conducted for the majority of the day.
- Shifters will provide support for these activities in the form of
- monitoring solenoid state
- controlled entry procedures
- other activities as requested
Data Collection¶
- 19:30
- Shifters and BLOC arrive and aid in preparation of the hall for data collection
- 20:00
- Start reference run (no more than 15 minutes)
- 20:30
- Setup beam for proposed half field run setting as a special run (settings to be provided by Jaroslaw)
- Run for 2000 pulses
- Setup beam for proposed half field run setting as a special run (settings to be provided by Jaroslaw)
- ~21:30
- Begin timing studies
- Settings at Ed Overton's discretion (probably reference run).
- Continue until end of shift.
Beamline Configurations¶
We expect to be running with 2V beam loss and beam bump active late in the day.
Reference run¶
Use the settings from run control
Half Field Beam line Test¶
To be provided by Jaroslaw- will have a pion beam assuming no DS.
Number of pulses | Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) | RunControl Tag |
# | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # | ||
2000 min | Pion | 178.25 | 173.18 | 169.42 | 29 | 37.35 | 68.20 | 41.69 | 130.66 | --- | 65.16 | 111.76 | 149.88 | 99.13 | 85.33 | 128.71 | 109.33 | Pion100_NoDS |
Run Time Preparedness.¶
- Downstream PID Detectors (EMR/KL/TOF)
- to be moved to final positions (July 16)
- HV to be turned on and DAQ reconfigured (July 16?)
- Detectors reconnected (July 15)
- Upstream PID Detectors (Ckov/TOFs)
- in final positions -- Done (July 4)
- Trackers
- Helium on
- downstream tracker -- Done but not maintained. (July 17)
- Upstream tracker still has a leak -- to be fixed in August
- Repair dead channels.
- Helium on
- Run Control
- CDB interface
- Need to improve stability of run control - remove need for IOC restart
- CDB beam line control
- Starting beam line current from zero doesn't work due to polarimeter -- Fixed (July 12)
- CDB interface
- Beam line control
- "Set all to 0" button has not operated in a predictable manner -- Corrected. (July 12)
- Beam line
- Decay Solenoid now filled (July 18)
- Ramp Sunday Afternoon?
- Correcting limits on the alarm handler.
- Likely to wait for September run.
- Decay Solenoid now filled (July 18)
Expected Work in the Hall¶
- Tracker commissioning and repairs to be conducted
- Will involve Paul R., Alan B. and Ed O.
- Expect them to be moving in and out of the hall
Notes¶
- Reminder: The end of run comment must be filled for the run to register as completed on CDB. Ensure that the enter key is struck while in the end comment dialogue box at the end of every run.
Temporary Run Control Instructions¶
- Restart Run Control after every run
- Currently using pro branch in MLCR.
- If Run Control was no restarted and it crashes the terminal, type "reset" (no quotes) to regain control, then restart Run Control. Do not log in/out
- If there is doubt contact Pierrick Hanlet or Chris Heidt
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Updated by Snopok, Pavel over 8 years ago ยท 10 revisions